The demand and supply function given as follow Qd= 100-10P and Qs= 10P
A. Find equilibrium price and quantity
B. Determine whether there is surplus or shortage at price equal to 6
C. What is the change in equilibrium price of demand curve change to Qd= 30- 10P, show the change graphically
1
Expert's answer
2021-09-07T07:15:01-0400

a)I)Equilibrium price


"100\u221210p=10p"


"100 = 20P"


"P =\frac{ 100}{20}"

​

"Equilibrium price (P) = 5"


Ii) Equilibrium quantity


"Qd = 100-10p"


"= 100-10(5)"


"= 100-50 = 50"


"Equilibrium quantity = 50"

b) The quantity supplied (QS) = "10P"


"= 10 \u00d7 6 = 60"


Indicating that there will be a surplus since it is more than the equilibrium quantity